# Collective Outlier Detection and Enumeration with Conformalized Closed Testing

### v2

@@ 4 Discussion
This section details how to empirically estimate the outlier distribution $G$, which is essential for implementing a data-driven version of Shirashi’s LMPI local testing procedure as described in Section 2.3.2. Recall from Section 2.3.2 that Shirashi’s optimality result (Shiraishi, 1985) is based on the following mixture model:
[…]
where $\theta\in[0,1]$ represents the proportion of outliers, and $G$ is a distribution function on $[0,1]$. In the following, we describe a practical and effective approach for computing an estimate $\hat{G}$ of $G$ using the data in $(X_{1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$, while ensuring that an approximately valid and approximately LMPI test can still be performed using Shirashi’s method conditional on $\hat{G}$.
@@
The key idea, inspired by approaches proposed in different contexts by Liang et al. (2024) and Marandon et al. (2024), is to randomly split the reference sample $(X_{1},\ldots,X_{m})$ into two disjoint subsets, namely $(X_{1},\ldots,X_{m_{1}})$ and $(X_{m_{1}+1},\ldots,X_{m})$, for some choice of $m_{1}$ such that $1<m_{1}<m$, like $m_{1}=\lceil m/2\rceil$. Intuitively, $(X_{1},\ldots,X_{m_{1}})$ is used as a smaller reference set and $(X_{m_{1}+1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$ as a “diluted” test set when fitting $G$. More precisely, the estimation of $G$ is carried out by fitting the following mixture model:
[…]
where $\theta_{1}=(\theta n+m_{1})/(n+m_{1})$ and $G$ is the same outlier distribution function as in Equation (A21). Any approach may be applied to fit this mixture model, as long as it is invariant to the ordering of the data in $(X_{m_{1}+1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$. A concrete approach is discussed in the next subsection.
The motivation for this method is that it ensures the inlier scores among $(Y_{1},\ldots,Y_{n})$ remain exchangeable with the smaller reference set $(X_{m_{1}+1},\ldots,X_{m})$ conditional on the estimated $\hat{G}$. Therefore, any downstream conformal inferences that utilizes $(X_{m_{1}+1},\ldots,X_{m})$ as a reference (or calibration) set are still valid. This is akin to the idea underlying AdaDetect, the individual outlier identification method proposed by Marandon et al. (2024).
A subtle point worth highlighting is that the asymptotic critical values for Shirashi’s LMPI test outlined in Section 2.3.2 were derived under the assumption that all inliers are i.i.d. (A21), whereas conditional on $\hat{G}$ the inliers among $(Y_{1},\ldots,Y_{n})$ are only exchangeable with $(X_{m_{1}+1},\ldots,X_{m})$ but not independent. However, to the best of our knowledge, this issue does not seem to have noticeable practical implications, even for only moderately large sample sizes. We refer to Section 2.6 for a more detailed discussion of this point.
@@
[…]
@@
While this approach may seem somewhat heuristic, as it is not aimed at making rigorous inferences about $G$, it is sufficient for our purposes. The invariance of this estimation procedure to the ordering of $(\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{m},\tilde{Y}_{1},\ldots,\tilde{Y}_{n})$ ensures the validity of our test regardless of how closely $\hat{G}$ approximates the true $G$. Moreover, as demonstrated in our numerical experiments, this estimation method performs well enough in practice to provide our testing method with a noticeable advantage over standard approaches, such as the WMW test, particularly in scenarios where the oracle LMPI test has a significant edge.
@@
When applied within a closed testing framework, Shirashi’s LMPI testing procedure can become computationally expensive without an appropriate shortcut. As discussed in Appendix A3, a fast and exact shortcut exists, but it relies on the monotonicity of the outlier probability density—specifically, the derivative of the function $\hat{G}$ estimated above. This motivates the need for an estimation procedure that guarantees a monotone outlier density.
@@
With Simes’ local test, the shortcut described in Goeman et al. (2019) and given in Algorithm 4 allows calculating $d_{\mathrm{Simes}}(S)$ in linear time, after an initial preparatory step that takes $\mathcal{O}(n\log n)$ time.
Input : $p$-values $p_{1},\ldots,p_{n}$; index set $S\subseteq[n]$; significance level $\alpha\in(0,1)$.
@@
4Compute $d_{\mathrm{Simes}}(S)=\min\{0\leq k\leq|S|:p_{(k+j:S)}>j\alpha/h\,\,\mathrm{for\,\,}j=1,\ldots,|S|-k\}$;
Output : A simultaneous $(1-\alpha)$-confidence lower bound $d_{\mathrm{Simes}}(S)$ for the number of outliers in $S$, for any $S\subseteq[n]$.
5
@@
Consequently, the two methods have the same weak FWER control and the same power for outlier detection. Furthermore, Goeman et al. (2019) showed that the confidence bound $d_{\mathrm{Simes}}$ is between the number of Hommel-FWER discoveries and the number of BH-FDR discoveries, i.e.
@@
The shortcut given in Algorithm 5 allows calculation of $d_{\mathrm{Storey-Simes}}(S)$ in quadratic time.
Input : $p$-values $p_{1},\ldots,p_{n}$; index set $S\subseteq[n]$; significance level $\alpha\in(0,1)$; tuning parameter $\lambda\in(0,1)$.
@@
2Initialize $d_{\mathrm{Storey-Simes}}(S)\leftarrow 0$;
3for each $i\in\{1,\ldots,|S|\}$ do
4       for each $j\in\{0,\ldots,n-|S|+i-1\}$ do
5

### v3

@@ 4 Discussion
The Edgington’s method has been studied by Rüschendorf (1982) and Meng (1994). They show that twice the average of the $p$-values is a valid $p$-value for arbitrary dependence. Recently, Choi and Kim (2023) demonstrated that the “twice the average” rule cannot be improved even under the assumption of exchangeability of the $p$-values. In contrast, the WMW test, when utilizing the average of the conformal $p$-values as the test statistic, rejects $H_{S}$ if $\bar{p}_{S}\leq\alpha^{\mathrm{Perm}}$, where the critical value $\alpha^{\mathrm{Perm}}$ is determined by taking the $\alpha$ quantile of the permutation distribution of $\bar{p}_{S}$, ensuring exact validity. This critical value $\alpha^{\mathrm{Perm}}$ may be calculated either exactly in small samples or through asymptotic approximations. A well-known asymptotic approach when $m$ and $|S|$ are both large is that of Hoeffding (1948), which leads to a much more powerful test compared to the combination approach for arbitrary dependence. In particular, the WMW test rejects $H_{S}$ when
$\bar{p}_{S}\leq$ $\,\frac{m+2}{2(m+1)}-z_{1-\alpha}\sqrt{\frac{m(|S|+m+1)}{12|S|(m+1)^{2}}},$
where $z_{1-\alpha}$ is the $(1-\alpha)$ quantile of the standard normal distribution. For $\alpha=0.1$ and $m=|S|=100$, this translates to rejecting $H_{S}$ if $\bar{p}_{S}\leq 0.453$, which is a much better critical value than the $\alpha/2=0.05$ required for arbitrary dependence.
Lehmann (1953) investigated the power of rank tests in the case when the distribution of the scores assigned to outliers in the test set is equal to $G(F)=F^{k}$ for any integer $k\geq 2$, which is referred to as Lehmann’s alternative (Balakrishnan, 2021). Here, $F^{k}$ represents the distribution of the maximum of $k$ independent random variables with distribution $F$. If $X\sim F$ and $Y\sim F^{k}$, then $\mathbb{P}(X<Y)=k/(k+1)$. It is worth noting that the distance of $F^{k}$ from $F$ remains the same for any specification of $F$.
Under Lehmann’s alternative $G(F)=F^{k}$, for integers $k\geq 2$, the associated density function is $g(u)=ku^{k-1}\mathds{1}[u\in[0,1]]$ and the test statistics in (4) reduces to:
$T^{\text{WMW, k}}$ $=\frac{k}{(N+1)\cdot\ldots\cdot(N+k-1)}\sum_{j=m+1}^{N}\prod_{l=0}^{k-2}(R_{j}+l),$ (A23)
which are the locally most powerful rank tests against the class of Lehmann’s alternatives. For $k=2$, the test statistic in (A23) differs from the WMW statistic only by a multiplicative constant and therefore yields an equivalent test.
Rank tests against Lehmann’s alternative were also studied in Conover and Salsburg (1988) and Rosenbaum (2007).
A1.5 Shirashi’s Locally Most Powerful Rank Test
Theorem A1 (Shiraishi (1985)).
Assume $X_{1},\ldots,X_{m}$ i.i.d. with continuous c.d.f. $F$, and $Y_{1},\ldots,Y_{n}$ i.i.d. with c.d.f. $(1-\theta)F+\theta G(F)$ for some $\theta\in[0,1]$ and some continuous c.d.f. $G$ on $[0,1]$ with density $g$. Then, the LMPR test rejects $H_{0}:\theta=0$ for large values of:
$T^{\text{g}}=\sum_{j=m+1}^{N}\mathbb{E}[g(U_{N}^{(R_{j})})].$
Under the setup of Section 2.3.2, we now write the local asymptotic power function of Shiraishi’s LMPR test and compare its relative efficiency with that of the Neyman–Pearson optimal test. Let $\pi^{\phi}_{N}(\theta)=\mathbb{E}_{\theta}(\phi)$ denote the power function of a test $\phi$ of size $\alpha$ for $\theta\in[0,1]$ and overall sample size $N$, and assume that $n/N\rightarrow\lambda\in(0,1)$. Consider a sequence of local alternatives $\theta_{N}=h/\sqrt{N}$ converging to the null hypothesis for some $h>0$. The local asymptotic power function of Shiraishi’s LMPR test $\phi^{g}$ is given by Shiraishi (1985, Theorem 1) as
$\pi^{\phi^{\mathrm{g}}}(h):=\lim_{N\rightarrow\infty}\pi^{\phi^{\mathrm{g}}}_{N}\Big(\frac{h}{\sqrt{N}}\Big)=\Phi\Big(z_{\alpha}+h\sqrt{I_{g}}\Big),\qquad I_{g}=\lambda(1-\lambda)\mathbb{V}\mathrm{ar}(g(U)),$
where $I_{g}$ is the local asymptotic information. The local asymptotic power is maximized at $\lambda=1/2$, so the Shiraishi’s test is most powerful for balanced calibration and test sample sizes. This follows from the fact that the number of cross-group comparisons is of order $mn=\lambda(1-\lambda)N^{2}$, which is maximized at $\lambda=1/2$. This observation provides a concrete and practically relevant guideline for test design.
By contrast, the local asymptotic power function of the optimal test $\phi^{opt}$, which is the likelihood ratio test (LRT) by the Neyman-Pearson lemma and assumes oracle knowledge of $F$, $G$ and $\theta$ (see also Appendix A4), is given by
$\pi^{\phi^{opt}}(h):=\lim_{N\rightarrow\infty}\pi^{\phi^{opt}}_{N}\Big(\frac{h}{\sqrt{N}}\Big)=\Phi\Big(z_{\alpha}+h\sqrt{I^{opt}_{g}}\Big),\qquad I^{opt}_{g}=\lambda\mathbb{V}\mathrm{ar}(g(U)).$
The power of the LRT increases with $\lambda$ because the test statistic depends only on the test sample. Since the inlier distribution $F$ is assumed known, the calibration sample is not used. It follows that the asymptotic relative efficiency (ARE), or Pitman efficiency, of Shiraishi’s LMPR test relative to the LRT optimal test equals $1-\lambda$ (Shiraishi, 1985, Corollary 1). This shortfall in asymptotic efficiency is not a limitation of the LMPR test per se, but rather a consequence of the two-sample setting. In the one-sample setting, the LMPR test may achieve optimal power, as shown in Shiraishi (1986).
The dense and sparse regimes $\sqrt{N}\theta_{N}\rightarrow\infty$ and $\sqrt{N}\theta_{N}\rightarrow 0$ are studied nonparametrically in Huang (2023) under a location-shift model for the outliers.
@@
This section details how to empirically estimate the outlier distribution $G$ and its density function $g$, which is essential for implementing a data-driven version of Shirashi’s local testing procedure as described in Section 2.3.2. Recall from Section 2.3.2 that Shirashi’s optimality result (Shiraishi, 1985) is based on the following mixture model:
$\begin{split}X_{1},\ldots,X_{m}&\overset{\text{i.i.d.}}{\sim}F,\\ Y_{1},\ldots,Y_{n}&\overset{\text{i.i.d.}}{\sim}(1-\theta)F+\theta G(F),\end{split}$ (A24)
where $\theta\in[0,1]$ represents the proportion of outliers, and $G$ is a distribution function on $[0,1]$ with density $g$. In the following, we describe a practical and effective approach for computing an estimate $\hat{g}$ of $g$ using the data in $(X_{1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$, while ensuring that an approximately valid and approximately LMPR test can still be performed using Shirashi’s method conditional on $\hat{g}$.
@@
The key idea, inspired by approaches proposed in different contexts by Liang et al. (2024) and Marandon et al. (2024), is to randomly split the calibration sample $(X_{1},\ldots,X_{m})$ into two disjoint subsets, namely $(X_{1},\ldots,X_{m_{1}})$ and $(X_{m_{1}+1},\ldots,X_{m})$, for some choice of $m_{1}$ such that $1<m_{1}<m$, like $m_{1}=\lceil m/2\rceil$. Intuitively, $(X_{1},\ldots,X_{m_{1}})$ is used as a smaller calibration set and $(X_{m_{1}+1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$ as a “diluted” test set when fitting $g$. More precisely, the estimation of $g$ is carried out by fitting the following mixture model:
$\begin{split}X_{1},\ldots,X_{m_{1}}&\overset{\text{i.i.d.}}{\sim}F,\\ X_{m_{1}+1},\ldots,X_{m},Y_{1},\ldots,Y_{n}&\overset{\text{i.i.d.}}{\sim}(1-\theta_{1})F+\theta_{1}G(F),\end{split}$ (A25)
where $\theta_{1}=(\theta n+m_{1})/(n+m_{1})$ and $G$ is the same outlier distribution function as in Equation (A24). Any approach may be applied to fit this mixture model, as long as it is invariant to the ordering of the data in $(X_{m_{1}+1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$. Otherwise, we simply need to permute the scores of the diluted test set before estimating $g$. The motivation for this method is that it ensures the inlier scores among $(Y_{1},\ldots,Y_{n})$ remain exchangeable with the smaller calibration set $(X_{m_{1}+1},\ldots,X_{m})$ conditional on the estimated $\hat{g}$. Therefore, any downstream conformal inferences that utilizes $(X_{m_{1}+1},\ldots,X_{m})$ as a reference (or calibration) set are still valid. This is akin to the idea underlying AdaDetect, the individual outlier identification method proposed by Marandon et al. (2024).
A subtle point worth highlighting is that conditional on $\hat{g}$ the inliers among $(Y_{1},\ldots,Y_{n})$ are only exchangeable with $(X_{m_{1}+1},\ldots,X_{m})$ but not independent.
Nevertheless, Shiraishi’s adaptive rank test, which replaces the unknown outlier density $g$ with its estimate $\hat{g}$, retains finite-sample validity, and the asymptotic approximation of its null distribution remains valid, as shown in Appendix A2.4.
@@
    $\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{m},\tilde{Y}_{1},\ldots,\tilde{Y}_{n}$ $\overset{\text{i.i.d.}}{\sim}(1-\theta_{1})\text{Uniform}(0,1)+\theta_{1}G,$ (A26)
@@
While this approach may seem somewhat heuristic, as it is not aimed at making rigorous inferences about $g$, it is sufficient for our purposes. The invariance of this estimation procedure to the ordering of $(\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{m},\tilde{Y}_{1},\ldots,\tilde{Y}_{n})$ ensures the validity of our test regardless of how closely $\hat{g}$ approximates the true $g$. Moreover, as demonstrated in our numerical experiments, this estimation method performs well enough in practice to provide our testing method with a noticeable advantage over standard approaches, such as the WMW test, particularly in scenarios where the oracle Shiraishi test has a significant edge.
@@
When applied within a closed testing framework, Shirashi’s testing procedure can become computationally expensive without an appropriate shortcut. As discussed in Appendix A3, a fast and exact shortcut exists, but it relies on the monotonicity of the outlier probability density
@@
A2.4 Shiraishi’s Adaptive Rank Test: Finite-Sample Validity and Asymptotic Null Distribution
Let $\hat{F}$ denote any estimator of $F$ based on $(X_{1},\ldots,X_{m_{1}})$. Define the probability integral-transformed variables $\tilde{X}_{j}=\hat{F}(X_{j})$, $j=m_{1}+1,\ldots,m$, and $\tilde{Y}_{j}=\hat{F}(Y_{j})$, $j\in[n]$, and estimate $g$ by fitting the mixture model
$\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{m},\tilde{Y}_{1},\ldots,\tilde{Y}_{n}\overset{\text{i.i.d.}}{\sim}(1-\theta_{1})\mathrm{Uniform}(0,1)+\theta_{1}G,$
where $G$ has density $g$ on $[0,1]$. Let $\hat{g}$ be the resulting estimator, and assume that it is invariant under permutations of $(\tilde{W}_{1},\ldots,\tilde{W}_{\tilde{N}})=(\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{m},\tilde{Y}_{1},\ldots,\tilde{Y}_{n})$, where $\tilde{N}=m-m_{1}+n$. No further assumptions on $\hat{g}$ are imposed beyond the condition (A29) discussed below, which ensures the validity of the asymptotic approximation. In particular, the i.i.d. assumption underlying the fitted mixture model need not hold.
Let $\tilde{R}_{1},\ldots,\tilde{R}_{\tilde{N}}$ denote the ranks of $\tilde{W}_{1},\ldots,\tilde{W}_{\tilde{N}}$. The Shiraishi adaptive rank test statistic is defined as
$T^{\hat{\text{g}}}$ $=\sum_{j=m-m_{1}+1}^{\tilde{N}}\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(\tilde{R}_{j})})],$ (A27)
and the corresponding test by
$\phi^{\hat{g}}=\mathds{1}\left\{T^{\hat{g}}>c_{\alpha}^{\hat{g}}\right\}$. (A28)
[…]
$\pi T^{\hat{g}}=\sum_{j=m-m_{1}+1}^{\tilde{N}}\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(\tilde{R}_{\pi(j)})})].$
The index $k$ is given by $k=\lceil(1-\alpha)N!\rceil$.
Finite-sample validity of the adaptive Shiraishi rank test follows from the exchangeability of the transformed vector $)(\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(\tilde{R}_{1})})],\ldots,\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(\tilde{R}_{\tilde{N}})}$)]), as shown formally in the following proposition.
Proposition A3.
If the vector $(\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{n},\tilde{Y}_{1},\ldots,\tilde{Y}_{n})$ is exchangeable, then the transformed vector
$)(\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(\tilde{R}_{1})})],\ldots,\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(\tilde{R}_{\tilde{N}})}$)])
is also exchangeable.
The proof relies on the following lemma, which states that a transformation of exchangeable random variables is itself exchangeable whenever applying a permutation after the transformation is equivalent to applying the transformation after permuting the original variables.
Lemma A1 (Dean and Verducci (1990)).
Suppose $(W_{1},\ldots,W_{k})\in\mathcal{W}^{k}$ is a vector of exchangeable random variables. Fix a transformation $A:\mathcal{W}^{k}\rightarrow(\mathcal{W}^{\prime})^{q}$. If for each permutation $\pi_{1}:[q]\rightarrow[q]$ there exists a permutation $\pi_{2}:[k]\rightarrow[k]$ such that
$\pi_{1}A(w)=A(\pi_{2}w)\quad\mathrm{for\,\,all\,\,}w\in\mathcal{W}^{k},$
then $A(\cdot)$ preserves exchangeability.
Proof of Proposition A3.
Let $A:\mathbb{R}^{\tilde{N}}\rightarrow\mathbb{R}^{\tilde{N}}$ denote the transformation that maps the vector $(\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{m},\tilde{Y}_{1},\ldots,\tilde{Y}_{n})$ to the vector $)(\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(\tilde{R}_{1})})],\ldots,\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(\tilde{R}_{\tilde{N}})}$)]). Let $\pi:[\tilde{N}]\rightarrow[\tilde{N}$] any permutation of the indices. Then for all $(w_{1},\ldots,w_{\tilde{N}})\in\mathbb{R}^{\tilde{N}}$
$).\pi A(w_{1},\ldots,w_{\tilde{N}})=A(\pi(w_{1},\ldots,w_{\tilde{N}}$)).
Thus, $A$ preserves exchangeability. ∎
We now state a parallel result to Theorem A2 for the asymptotic approximation of the null distribution of Shiraishi’s adaptive test.
Theorem A2.
Suppose that
$\frac{\max_{1\leq r\leq\tilde{N}}(\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(r)})]-\hat{\mu}_{\tilde{N}})^{2}}{\min(n,m-m_{1})\hat{\sigma}^{2}_{\tilde{N}}}\xrightarrow{\mathbb{P}}0,$ (A29)
where
$\hat{\mu}_{\tilde{N}}=\frac{1}{\tilde{N}}\sum_{r\in[\tilde{N}]}\mathbb{E}[\hat{g}(U_{N}^{(r)})],\qquad\hat{\sigma}^{2}_{\tilde{N}}=\frac{1}{\tilde{N}-1}\sum_{r\in[\tilde{N}]}(\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(r)})]-\hat{\mu}_{\tilde{N}})^{2},$
and where $\hat{g}$ is invariant under permutations of the pooled sample $(\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{m},\tilde{Y}_{1},\ldots,\tilde{Y}_{n})$. Then, under the null hypothesis $\tilde{H}^{*}_{0}:\mathrm{the\,\,vector\,\,}(\tilde{X}_{m_{1}+1},\ldots,\tilde{X}_{m},\tilde{Y}_{1},\ldots,\tilde{Y}_{n})\,\,\mathrm{is\,\,exchangeable},$ the standardized Shiraishi adaptive statistic converges in distribution to a standard normal,
$\frac{T^{\hat{\mathrm{g}}}-n\hat{\mu}_{N}}{\sqrt{\frac{(m-m_{1})n}{\tilde{N}}\hat{\sigma}^{2}_{\tilde{N}}}}\xrightarrow{d}N(0,1).$ (A30)
Proof.
[…]
The Shiraishi adaptive statistic can be written as
$T^{\hat{\text{g}}}$ $=\sum_{r\in[\tilde{N}]}\mathbb{E}[\hat{g}(U_{\tilde{N}}^{(r)})]C_{r},$
[…]
Note that if $\hat{g}$ is bounded in probability, then condition (A29) holds. This property is satisfied, for example, by the nonparametric estimator $\hat{g}$ of Patra and Sen (2016), and may alternatively be ensured via truncation of $\hat{g}$.
A2.5 Shiraishi’s Adaptive Rank Test: Local Asymptotic Power
We now show that, under suitable assumptions, the estimator $\hat{g}$ of Patra and Sen (2016) satisfies condition (9). Consequently, the adaptive rank test based on the Patra-Sen estimator has the same local asymptotic distribution as the oracle LMPR test and therefore achieves the same local asymptotic power. Moreover, the closed-testing shortcut described in Appendix A3.4 applies, giving an adaptive test that is directly usable in practice.
In the procedure described in Appendix A2, the calibration sample $(X_{1},\ldots,X_{m})$ is split into two parts. We estimate $F$ by the empirical distribution function $\hat{F}$ based on $X_{1},\ldots,X_{m_{1}}$. Define the probability integral-transformed variables $\tilde{X}_{j}=\hat{F}(X_{j})$, $j=m_{1}+1,\ldots,m$, and $\tilde{Y}_{j}=\hat{F}(Y_{j})$, $j\in[n]$, so that the inlier distribution is discrete Uniform on $[0,1]$. The distribution of the inlier can be made exactly Uniform on $[0,1]$ by applying the randomized probability integral-transform
$\tilde{W}=\hat{F}(W-\epsilon)+U[\hat{F}(W)-\hat{F}(W-\epsilon)]$
for a sufficiently small $\epsilon>0$, where $U\sim\mathrm{Uniform}(0,1)$.
Let $\tilde{W}_{1},\ldots,\tilde{W}_{\tilde{N}}$ denote the pooled transformed sample, where $\tilde{N}=(m-m_{1})+n$. Conditional on $(X_{1},\ldots,X_{m_{1}})$, the pooled sample satisfies
$\tilde{W}_{1},\ldots,\tilde{W}_{\tilde{N}}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}(1-\theta_{1})\mathrm{Uniform}(0,1)+\theta_{1}G.$
We further assume that $G$ has a continuous, bounded, and non-increasing density $g$, and the mixture model is identifiable, for example by the condition $g(1)=0$, placing us in the setting required for the estimator $\hat{g}_{\tilde{N}}$ of Patra and Sen (2016).
The Grenander-type estimator $\hat{g}_{\tilde{N}}$ proposed by Patra and Sen (2016) satisfies, for any $p\in[1,2)$,
$\int_{0}^{1}\bigl|\hat{g}_{\tilde{N}}(u)-g(u)\bigr|^{p}\,du=O_{\mathbb{P}}\!\left(\tilde{N}^{-p/3}\right),$
provided that assumptions (A1), (A2), and (A2) in Durot (2007) are satisfied; see Corollary 1 therein.
The estimation error $(\hat{g}_{\tilde{N}}-g)$ is uniformly bounded in probability because both terms are uniformly bounded. This boundedness, together with the above $L^{p}$ convergence, implies convergence in $L^{2}$: for any $p\in[1,2)$,
$\int_{0}^{1}\bigl|\hat{g}_{\tilde{N}}(u)-g(u)\bigr|^{2}\,du\leq\biggl(\sup_{u\in[0,1]}\bigl|\hat{g}_{\tilde{N}}(u)-g(u)\bigr|\biggr)^{2-p}\int_{0}^{1}\bigl|\hat{g}_{\tilde{N}}(u)-g(u)\bigr|^{p}\,du\xrightarrow{\mathbb{P}}0.$
Let $u_{r}=r/(\tilde{N}+1)$ and set $\delta_{\tilde{N}}(u)=\hat{g}_{\tilde{N}}(u)-g(u)$. Then
$\frac{1}{\tilde{N}}\sum_{r=1}^{\tilde{N}}(\hat{b}_{r}-b_{r})^{2}=\int_{0}^{1}\delta_{\tilde{N}}(u)^{2}\,du+E_{\tilde{N}}$,
where
$E_{\tilde{N}}:=\frac{1}{\tilde{N}}\sum_{r=1}^{\tilde{N}}\delta_{\tilde{N}}(u_{r})^{2}-\int_{0}^{1}\delta_{\tilde{N}}(u)^{2}\,du.$
[…]
$\mathrm{TV}\!\left(\delta_{\tilde{N}}^{2}\right)\leq 2\|\delta_{\tilde{N}}\|_{\infty}\,\mathrm{TV}(\delta_{\tilde{N}})=O_{\mathbb{P}}(1).$
It follows that
$|E_{\tilde{N}}|\leq\frac{\mathrm{TV}(\delta_{\tilde{N}}^{2})}{\tilde{N}}=o_{\mathbb{P}}(1),$
and therefore
$\frac{1}{\tilde{N}}\sum_{r=1}^{\tilde{N}}(\hat{b}_{r}-b_{r})^{2}\xrightarrow{\mathbb{P}}0.$
Other nonparametric estimators for the unknown outlier density are studied in Matias and others (2014) and Liang and Kou (2025).
@@
Shortcuts fall into two categories: exact shortcuts that yield identical rejections to the full closed testing procedure, and approximate or conservative shortcuts that may produce fewer rejections than the full procedure. The shortcuts presented in this section are exact, with the exception of an approximate shortcut discussed at the end of the section, based on the method of Gazin et al. (2024).
The table below summarizes the computational complexity of these shortcuts for computing the lower bound $d$ on the total number of outliers, as well as the lower bound $d(S)$ for a subset $S$ of the test sample. In all exact shortcuts considered in this paper, computation of $d(S)$ requires first computing $d$. Since $d$ does not depend on $S$, it needs to be computed only once.
The reported computational complexity includes a preparatory step in which the $m$ calibration and $n$ test scores are sorted, the ranks of the pooled scores are computed, and the conformal $p$-values are obtained from these ranks, and are therefore already ordered. This requires $\mathcal{O}(n\log n)$ time when $m$ is of the same order as $n$. Existing shortcuts for Simes, WMW, and Fisher local tests are implemented efficiently, with computational complexity linearithmic in $n$ for both $d$ and $d(S)$, matching the cost of sorting $n$ values. The quadratic and cubic complexities for computing $d$ and $d(S)$ for Storey-Simes and Shiraishi local tests arise from naive implementations of Algorithms 5 and 6, and could likely be further refined with additional effort.
Local test Computational complexity for computing the lower bound
for the total number of outliers for the number of outliers in a subset
Simes $\mathcal{O}(n\log n)$ $\mathcal{O}(n\log n)$
Storey-Simes $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{3})$
Shiraishi $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{3})$
WMW / Fisher $\mathcal{O}(n\log n)$ $\mathcal{O}(n\log n)$
Table A3: Computational cost of the shortcut procedures for Algorithm 1 for computing the lower bound $d$ for the total number of outliers and the lower bound $d(S)$ on the number of outliers in a subset $S$ of the test set, as a function of the chosen local test.
@@
With Simes’ local test, the shortcut described in Goeman et al. (2019) and given in Algorithm 4 is exact and allows calculating $d_{\mathrm{Simes}}(S)$ in linear time, after an initial preparatory step that takes $\mathcal{O}(n\log n)$ time.
Input : $p$-values $p_{1},\ldots,p_{n}$; significance level $\alpha\in(0,1)$.
@@
4Compute the lower bound on the total number of outliers $d_{\mathrm{Simes}}=d_{\mathrm{Simes}}([n])=n-h$;
5Query $S\subset[n]$: compute $d_{\mathrm{Simes}}(S)=\min\{0\leq k\leq|S|:p_{(k+j:S)}>j\alpha/h\,\,\mathrm{for\,\,}j=1,\ldots,|S|-k\}$;
Output : A simultaneous $(1-\alpha)$-confidence lower bound $d_{\mathrm{Simes}}(S)$ for the number of outliers in $S$, for $S=[n]$ and for any queried subset $S\subset[n]$.
6
@@
The first relation, $1\{\phi_{S}=1\}\geq 1\{d(S)>0\}$, follows from the identity $1\{d(S)>0\}=1\{\phi_{K}=1,\forall\,K\supseteq S\}$. It highlights that a procedure which fixes $S$ a priori and tests only $H_{S}$ is at least as powerful as testing the same hypothesis with correction for multiple testing. In the relevant case $S=[n]$, however, the two approaches coincide. The second relation emphasizes that the number of outliers obtained by enumeration is at least as large as the number of outliers identified.
@@
Consequently, the two methods have the same weak FWER control and the same power for outlier detection.
This is illustrated in the right panel of Figure 1, where the solid gray line (Simes) and the dotted gray line (BH) coincide, indicating that both procedures have identical power for rejecting the global null. Furthermore, Goeman et al. (2019) showed that the confidence bound $d_{\mathrm{Simes}}$ is between the number of Hommel-FWER discoveries and the number of BH-FDR discoveries, i.e.
@@
The shortcut given in Algorithm 5 allows calculation of $d_{\mathrm{Storey-Simes}}(S)$ in cubic time. Algorithm 5 is based on a worst-case argument. For any $L\subseteq[n]$ with $|L|=l$, the Storey-Simes test rejects $H_{L}$ if
$\min_{j\in L}\Big(l\frac{p_{(j:L)}}{j}\Big)\leq\alpha\Big(\frac{l(1-\lambda)}{1+\sum_{j\in[L]}\mathds{1}\{p_{j}>\lambda\}}\Big).$
Among all subsets $L$ of size $l$, the least favorable (hardest to reject) choice is the set corresponding to the largest $l$ $p$-values $p_{(n-l+1)},\ldots,p_{(n)}$. Consequently, for a fixed $l$, it suffices to evaluate the Storey–Simes test on this worst-case subset. Algorithm 5 exploits this fact to implement closed testing exactly, while avoiding explicit enumeration of all subsets.
Input : $p$-values $p_{1},\ldots,p_{n}$; significance level $\alpha\in(0,1)$; tuning parameter $\lambda\in(0,1)$.
@@
2Sort the $p$-values $p_{(1)}\leq\ldots\leq p_{(n)}$;
[…]
4while $h\geq 1$ and $\min_{j\in[h]}\Big(h\frac{p_{(n-h+j)}}{j}\Big)\leq\alpha\Big(\frac{h(1-\lambda)}{1+\sum_{j\in[h]}\mathds{1}\{p_{(n-h+j)}>\lambda\}}\Big)$ do
5 $h\leftarrow h-1$;
6
7Compute $d_{\mathrm{Storey-Simes}}=d_{\mathrm{Storey-Simes}}([n])=n-h$;
8Query $S\subset[n]$: $d_{\mathrm{Storey-Simes}}(S)\leftarrow 0$;
9Sort the $p$-values in $S$, $p_{(1:S)}\leq\ldots\leq p_{(|S|:S)}$, and in $S^{c}$, $p_{(1:S^{c})}\leq\ldots\leq p_{(n-|S|:S^{c})}$;
10for each $i\in\{1,\ldots,|S|\}$ do
11 for each $j\in\{0,\ldots,\max(h-|S|+i-1,0)\}$ do
12